OpenStudy (anonymous):

The formula for the amount with simple interest is A = P(1 + rt). Find r, if A = 224, P = 200, and t = 2. r = 0.6 r = 0.12 r = 0.06 r = 0.012

OpenStudy (jiteshmeghwal9):

A=p(1+rt) 224=200(1+r2) 224=200+200r2 224-200=200r2 4=200r2 4/200=r2 1/50=r2 1/50 times 2=r 1/25=r